-- Details --
- The lock duration was about 1 hour. I kicked the IFO for measurement, so the lock was stable enough to continue to be locked for several hours.
- The guardian can hand-off the TRAS signal to the OMC DC signal, although the OMC cannot be locked automatically. So the LSC LOCK guardian waits for the OMC to be locked after putting the DARM offset. I took over the OMC operation to Yokozawa-san and Yamamoto-sensei.
- The noise limiting the sensitivity will be reported by Yamamoto-sensei.
- DC offset is 30 pm, corresponds to 20 mW (2000cnts at OMC DC PD A). I put -0.5 cnts to TRAS servo filter.
- OMC UGF is ~100 Hz. The gain of FB_FLT is 0.5 and FM2, FM6 was engaged. Once the OMC is locked, FM5 was engaged.
- Dr. Okutomi made a beautiful IFO monitor screen as shown in the attached picture. We can easily find the BNS range. I really like it.
The OMC DC PD B has a noise around 1 kHz, which is not seen in PD A. Is this known?